# Accelerating Training Speed of Tiny Recursive Models via Curriculum-Guided Adaptive Recursion

### v1

@@ I Introduction
The recent surge in large language models (LLMs) with hundreds of billions of parameters has demonstrated remarkable capabilities across diverse tasks [brown2020gpt3, chowdhery2023palm]. However, this paradigm of scaling model size incurs prohibitive computational costs during both training and inference, limiting accessibility and deployability. An alternative direction has emerged through test-time computation scaling [snell2024scaling, zhang2024encode], where smaller models iteratively refine their outputs through multiple reasoning steps, trading inference cycles for model parameters.
Building upon this principle, the recently proposed Hierarchical Reasoning Model (HRM) [chu2024hierarchical] and its simplified variant, Tiny Recursive Model (TRM) [jolicoeur2024trm], have shown that networks with merely 7M parameters can match or exceed LLMs ten thousand times their size on hard reasoning tasks such as Sudoku, maze solving, and ARC-AGI [chollet2019measure]. The key insight is recursive reasoning: a tiny network iteratively refines solutions through nested recursion cycles and deep supervision [lee2015deeply], effectively emulating deep architectures while maintaining parameter efficiency. Through adaptive computation time [graves2016adaptive], TRM achieves strong performance on reasoning benchmarks using only 7M parameters. However, training these models remains computationally expensive, with the original TRM paper reporting approximately 36 GPU-hours for Sudoku extreme dataset, limiting broader adoption and rapid experimentation.
@@
We propose CGAR (Curriculum-Guided Adaptive Recursion), a novel training methodology that fundamentally rethinks how recursive models learn. Unlike all prior curriculum learning approaches that focus on data ordering [bengio2009curriculum, hacohen2019power, soviany2022curriculum], sample weighting, or parameter reduction [hu2022lora], CGAR introduces the first curriculum on architectural recursion depth itself. This paradigm shift recognizes that the effective computational depth $\mathcal{D}_{\text{eff}}(n,T)$ of recursive architectures is a learnable training hyperparameter that should adapt with optimization progress, rather than remain fixed throughout training. CGAR introduces two synergistic innovations. First, Progressive Depth Curriculum (PDC) dynamically schedules recursion parameters $(n,T)$ based on normalized training progress $\rho=e/E$, implementing a three-stage curriculum that transitions from shallow $(2,1)$ through medium $(4,2)$ to full depth $(6,3)$ configurations, providing 41.4% FLOPs reduction while preventing early-stage overfitting. Second, Hierarchical Supervision Weighting (HSW) applies principled exponential decay $w_{t}=\lambda^{t-1}/Z_{\lambda}$ to supervision steps, derived from empirical observations of gradient magnitude decay in recursive architectures, achieving 40% gradient variance reduction and accelerated convergence through improved signal-to-noise ratio in stochastic gradients.
@@
The remainder of this paper is organized as follows: Section II reviews related work. Section III presents the CGAR framework. Section IV describes experimental setup. Section V reports results and ablations. Section VI concludes with limitations and future directions.
@@ II Related Work
Curriculum learning [bengio2009curriculum] proposes training neural networks by gradually increasing data difficulty, ordering samples from simple to complex to improve convergence. Subsequent work [graves2017automated, soviany2022curriculum, zhou2023curriculum] has expanded this paradigm through automated difficulty adjustment and applications across diverse domains. Progressive architectures [rusu2016progressive, huang2016stochastic] adjust network capacity during training, while recent approaches apply curriculum to neural architecture search [zhou2022curriculum, guo2020breaking]. Tang et al. [tang2023progressive] recently explored progressive learning from shallow to deep representations in IEEE TNNLS, demonstrating improved generalization through staged depth increases, though their approach focuses on layer-wise pretraining rather than dynamic recursion scheduling. Complementary approaches address training efficiency through architectural innovations: Tan & Le [tan2021efficientnetv2] combine progressive image sizing with neural architecture search; Hu et al. [hu2022lora] achieve parameter efficiency via low-rank adaptation; Zhuang et al. [zhuang2023efficient, tay2022efficient] survey efficient training methods for transformers. Cao & Tsang [cao2024adaptive] propose adaptive momentum for DNNs training in IEEE TNNLS, eliminating momentum hyperparameter tuning and accelerating convergence, achieving significant error reduction on CIFAR datasets. However, all existing curriculum methods focus on data ordering, sample weighting, or parameter reduction, treating architectural depth as fixed.
Adaptive computation mechanisms enable dynamic resource allocation based on input complexity. Graves [graves2016adaptive] introduces Adaptive Computation Time (ACT) for RNNs with learned halting at inference time. PonderNet [banino2021pondernet] learns variable computation steps, while SkipNet [wang2018skipnet] and BERxiT [xin2021berxit] enable early exiting in convolutional and transformer networks. Recent work on Mixture-of-Depths [raposo2024mixture] dynamically allocates compute in language models. Han et al. [han2021dynamic] survey dynamic networks that adjust depth, width, or resolution at inference, while Snell et al. [snell2024scaling] demonstrate test-time computation scaling for reasoning tasks. Deep supervision [lee2015deeply] addresses training efficiency by attaching auxiliary losses to intermediate layers. Chen et al. [chen2018gradnorm] propose GradNorm for dynamic loss weight balancing via gradient magnitude equalization, while Zhang et al. [zhang2024hierarchical] explore hierarchical supervision strategies for multi-task learning in IEEE TNNLS, though their weighting focuses on task relationships rather than temporal decay in recursive architectures. TRM [jolicoeur2024trm] applies uniform supervision weighting across reasoning steps, though recursive architectures exhibit exponentially diminishing gradients at later steps due to cumulative detachment operations. Recent recursive architectures revisit iterative refinement for complex reasoning. Universal Transformers [dehghani2019universal] apply recurrent mechanisms to transformer blocks for iterative processing. Self-refinement approaches [madaan2023selfrefine, shinn2023reflexion] enable models to improve outputs through self-feedback and verbal reinforcement, while THREAD [sun2024thread] explores recursive spawning for deeper reasoning. Qasim et al. [qasim2025recursive] propose recursive decomposition of logical thoughts for superior reasoning in large language models. Chen et al. [chen2018neural] introduce Neural ODEs parameterizing continuous-depth networks; Wang et al. [wang2024loopresidual] propose Loop-Residual Networks. Most relevant, Chu et al. [chu2024hierarchical] develop the Hierarchical Reasoning Model (HRM) with dual-latent recursion requiring $\sim\!\!72$ GPU-hours training. Jolicoeur-Martineau [jolicoeur2024trm] simplifies HRM into TRM, demonstrating $7\text{M}$ parameter models match LLMs $10{,}000\times$ larger on reasoning benchmarks, though training still requires $\sim\!\!36$ GPU-hours.
@@
CGAR differs fundamentally from prior work. Unlike data-based curricula and parameter-efficient methods, we introduce curriculum learning on recursion depth $(n,T)$, dynamically adjusting effective computational layers $\mathcal{D}_{\text{eff}}(n,T)$ during training based on epoch progress. Unlike ACT’s inference-time adaptation, our progressive depth curriculum operates at training time via deterministic scheduling. Unlike uniform supervision, our hierarchical weighting $w_{t}=\lambda^{t-1}$ accounts for gradient magnitude decay specific to recursive architectures. CGAR is a training methodology applicable to any recursive model, achieving $1.71\times$ speedup (10.9310.93h $\to$ 6.386.38h on A100) while maintaining competitive accuracy. Ablations reveal progressive depth curriculum alone provides $2.26\times$ speedup with improved accuracy, demonstrating architectural curriculum benefits beyond computational savings.
@@ III Method
Our goal is to accelerate the training speed of Tiny Recursive Model (TRM) [jolicoeur2024trm], which suffers from training inefficiency due to fixed-depth training at all epochs (the original TRM paper reports $\sim\!36$ GPU-hours per dataset on their hardware). We propose CGAR (Curriculum-Guided Adaptive Recursion), a training methodology that achieves $1.71\times$ speedup through two complementary techniques: progressive depth curriculum (PDC) and hierarchical supervision weighting (HSW). Under controlled conditions on A100 GPU, CGAR reduces training time from $10.93$ hours (our replicated TRM baseline) to $6.38$ hours. We begin by reviewing TRM’s architecture and identifying its training inefficiencies, then present our solutions with mathematical formulations and theoretical analysis.
@@ III-A Background: TRM Architecture and Training
TRM trains through deep supervision [lee2015deeply], attaching losses to all $N_{\text{sup}}$ steps with uniform weighting $w_{t}=1/N_{\text{sup}}$. Let $h_{\text{out}}:\mathbb{R}^{L\times D}\rightarrow\mathbb{R}^{L\times V}$ denote the output projection head. The training objective is:
@@
Figure 2: Algorithm 1. CGAR Training with Progressive Curriculum and Hierarchical Weighting
@@ IV Experiments
Following TRM [jolicoeur2024trm], we report exact match accuracy $\text{Acc}_{\text{exact}}:=N^{-1}\sum_{i=1}^{N}\mathds{1}[\hat{\bm{y}}_{i}=\bm{y}_{i}^{*}]$ measuring puzzles with all $81$ tokens correct (primary metric, since partial solutions violate global constraints), and token accuracy $\text{Acc}_{\text{token}}:=(NL)^{-1}\sum_{i=1}^{N}\sum_{j=1}^{L}\mathds{1}[\hat{y}_{ij}=y_{ij}^{*}]$ measuring per-position correctness (auxiliary metric). We evaluate CGAR on Sudoku-Extreme to answer four key research questions:
@@ IV-A Model Configuration
We implement CGAR using TinyRecursiveReasoningModel architecture with hidden dimension $D=512$, transformer blocks having $n_{L}=2$ layers with $h=8$ attention heads and feed-forward dimension $D_{\text{FFN}}=2048$, recursion depth $(n,T)$ following curriculum $\mathcal{C}_{\text{PDC}}(\rho)$ from Section III, deep supervision at $N_{\text{sup}}=16$ steps with hierarchical weighting $w_{t}=\lambda^{t-1}/Z_{\lambda}$ where $\lambda=0.7$ and $Z_{0.7}=3.283$, and total parameters $p\approx 5.0\text{M}$ identical to baseline TRM. The three-stage curriculum transitions at progress $\rho\in\{0.3,0.6,1.0\}$ with depths $(n,T)=(2,1)$ giving $\mathcal{D}_{\text{eff}}=6$ for $\rho<0.3$, then $(4,2)$ with $\mathcal{D}_{\text{eff}}=20$ for $0.3\leq\rho<0.6$, and $(6,3)$ with $\mathcal{D}_{\text{eff}}=42$ for $\rho\geq 0.6$. For controlled comparison, we train a baseline TRM on identical hardware (A100 GPU) using fixed $(n,T)=(6,3)$ with uniform weighting $w_{t}=1/16$ across all epochs following the TRM protocol [jolicoeur2024trm].
@@ IV-B Training Configuration
We optimize with AdamW [loshchilov2017adamw] using learning rate $\eta=5\times 10^{-4}$ with cosine annealing linear warmup [ma2024why] over 1,0001{,}000 steps ($\sim\!1.5\%$ of training), weight decay $\lambda_{\text{wd}}=0.01$, batch size $B=768$ on single GPU, gradient clipping at max norm $1.0$, and FP16 mixed precision with automatic loss scaling. Training runs for E=50, [^0] E=50{,}000 epochs following TRM protocol, where each epoch processes one minibatch of $B$ samples yielding $\sim\!65{,}000$ total optimization steps. We save checkpoints every 5,0005{,}000 epochs for progressive evaluation. All experiments use single NVIDIA A100 GPU (80GB VRAM) with CUDA 11.8 and PyTorch 2.0. Under these controlled conditions, CGAR training completes in $6.38$ hours wall-clock time versus $10.93$ hours for our replicated baseline TRM, achieving $1.71\times$ speedup.
@@ IV-C Implementation Details
Following TRM [jolicoeur2024trm], we detach gradients for first T1T-1 H-cycles (Algorithm 2 lines 11-15), backpropagating only through final cycle $j=T$, preventing gradient explosion while maintaining supervision at all $N_{\text{sup}}$ steps. We implement ACT [graves2016adaptive] with learned halting head $h_{\text{halt}}:\mathbb{R}^{L\times D}\rightarrow[0,1]$ predicting probabilities $q_{t}\in[0,1]$ supervised via binary cross-entropy $\ell_{\text{BCE}}(q_{t},\mathds{1}[\hat{\bm{y}}_{t}=\bm{y}^{*}])$ during training, halting at inference when $q_{t}>0.5$ for all positions. Curriculum progress $\rho_{e}=e/E$ computes at epoch start, determining depth $(n_{e},T_{e})=\mathcal{C}_{\text{PDC}}(\rho_{e})$ applied uniformly across all samples that epoch, ensuring stable gradient statistics within optimization steps. For reproducibility, we set random seed $42$ for PyTorch and NumPy.
@@ IV-D Evaluation Protocol
During training, we log primary loss $\mathcal{L}_{\text{CGAR}}$ from Eq. (6), training exact and token accuracies on current minibatch, curriculum progress $\rho$ with current depth $(n,T)$, ACT halting accuracy and loss, and gradient norms $\|\nabla_{\theta}\mathcal{L}\|_{2}$ every $100$ optimization steps. After training completion, we evaluate all $5$ saved checkpoints (epochs 30,00030{,}000 to 50,00050{,}000 at 5,0005{,}000 intervals) on full held-out test set ($N_{\text{test}}=200{,}000$ puzzles), computing exact and token accuracies, measuring inference speed (puzzles/second), analyzing halting step distributions $\{\delta(\bm{x}):\bm{x}\in\mathcal{D}_{\text{test}}\}$, and calculating generalization gap $\mathcal{R}(\theta):=\text{Acc}_{\text{test}}-\text{Acc}_{\text{train}}$. We compare CGAR against our replicated TRM baseline trained on identical hardware (A100 GPU) under the same experimental conditions. Note that the original TRM paper [jolicoeur2024trm] reports $87.4\%$ exact accuracy with $\sim\!36$ hours training on different hardware, but we conduct our primary comparison using our controlled baseline for fairness.

### v2

# Accelerating Training of Recursive Reasoning Models with Curriculum Guided Adaptive Recursion

@@ I Introduction
The recent surge in large language models (LLMs) with hundreds of billions of parameters has demonstrated remarkable capabilities across diverse tasks [1, 2]. However, this paradigm of scaling model size incurs prohibitive computational costs during both training and inference, limiting accessibility and deployability. An alternative direction has emerged through test-time computation scaling [3, 4], where smaller models iteratively refine their outputs through multiple reasoning steps, trading inference cycles for model parameters.
Building upon this principle, the recently proposed Hierarchical Reasoning Model (HRM) [5] and its simplified variant, Tiny Recursive Model (TRM) [6], have shown that networks with merely 7M parameters can match or exceed LLMs ten thousand times their size on hard reasoning tasks such as Sudoku, maze solving, and ARC-AGI [7]. The key insight is recursive reasoning: a tiny network iteratively refines solutions through nested recursion cycles and deep supervision [8], effectively emulating deep architectures while maintaining parameter efficiency. Through adaptive computation time [9], TRM achieves strong performance on reasoning benchmarks using only 7M parameters. However, training these models remains computationally expensive, with the original TRM paper reporting approximately 36 GPU-hours for Sudoku extreme dataset, limiting broader adoption and rapid experimentation.
@@
We propose CGAR (Curriculum-Guided Adaptive Recursion), a novel training methodology that fundamentally rethinks how recursive models learn. Unlike all prior curriculum learning approaches that focus on data ordering [10, 11, 12], sample weighting, or parameter reduction [13], CGAR introduces the first curriculum on architectural recursion depth itself. This paradigm shift recognizes that the effective computational depth $\mathcal{D}_{\text{eff}}(n,T)$ of recursive architectures is a learnable training hyperparameter that should adapt with optimization progress, rather than remain fixed throughout training. CGAR introduces two synergistic innovations. First, Progressive Depth Curriculum (PDC) dynamically schedules recursion parameters $(n,T)$ based on normalized training progress $\rho=e/E$, implementing a three-stage curriculum that transitions from shallow $(2,1)$ through medium $(4,2)$ to full depth $(6,3)$ configurations, providing 41.4% FLOPs reduction while preventing early-stage overfitting. Second, Hierarchical Supervision Weighting (HSW) applies principled exponential decay $w_{t}=\lambda^{t-1}/Z_{\lambda}$ to supervision steps, derived from empirical observations of gradient magnitude decay in recursive architectures, achieving 40% gradient variance reduction and accelerated convergence through improved signal-to-noise ratio in stochastic gradients.
@@
The remainder of this paper is organized as follows: Section II reviews related work. Section III presents the CGAR framework. Section IV describes experimental setup. Section V reports results and ablations. Section VI concludes with limitations and future directions. The source code and trained models will be made publicly available upon publication.
@@ II Related Work
Curriculum learning [10] trains neural networks on progressively difficult examples, improving convergence and generalization across diverse domains [14, 12, 15]. Progressive architectures adjust network capacity during training: Rusu et al. [16] incrementally expand capacity for continual learning, while Huang et al. [17] introduced stochastic depth to vary effective network depth. Recent work applies curriculum to neural architecture search [18, 19]. Tang et al. [20] investigated progressive learning from shallow to deep representations via layer-wise pretraining. Training efficiency has been addressed through architectural innovations [21, 13, 22, 23]. Cao and Tsang [24] proposed adaptive momentum for training, eliminating hyperparameter tuning. However, existing curriculum methods focus on data ordering, sample weighting, or parameter reduction, treating architectural depth as fixed.
Adaptive computation mechanisms dynamically allocate resources based on input complexity. Graves [9] introduced Adaptive Computation Time (ACT) for RNNs with learned halting at inference. PonderNet [25] learns variable computation steps, while early exit strategies reduce inference costs in CNNs [26] and transformers [27]. Raposo et al. [28] proposed Mixture-of-Depths for language models. Han et al. [29] survey dynamic networks adjusting depth, width, or resolution at inference, while Snell et al. [3] demonstrate test-time computation scaling for reasoning. Deep supervision [8] attaches auxiliary losses to intermediate layers for effective gradient flow. Chen et al. [30] introduced GradNorm for dynamic loss balancing, and Zhang et al. [31] investigated hierarchical supervision for multi-task learning. These approaches do not account for temporal decay in recursive architectures.
Recursive reasoning mechanisms have gained attention for complex problem solving. Universal Transformers [32] apply recurrent processing to transformer blocks. Self-refinement methods [33, 34] enable iterative improvement through self-feedback. Sun et al. [35] explored recursive spawning, while Qasim et al. [36] developed recursive decomposition for enhanced LLM reasoning. Chen et al. [37] introduced Neural ODEs for continuous-depth networks, and Wang et al. [38] proposed Loop-Residual Networks. Most relevant, Chu et al. [5] developed the Hierarchical Reasoning Model (HRM) requiring approximately 72 GPU-hours training. Jolicoeur-Martineau [6] simplified HRM into TRM, demonstrating 7M parameter models matching LLMs $10{,}000\times$ larger, though training requires approximately 36 GPU-hours.
@@
CGAR differs fundamentally from prior work. Unlike data-based curricula [10, 11, 12] and parameter-efficient methods [13], we introduce curriculum learning on architectural recursion depth $(n,T)$, dynamically adjusting effective computational depth $\mathcal{D}_{\text{eff}}(n,T)$ during training. Unlike inference-time adaptation [9, 25], our progressive depth curriculum operates at training time via deterministic scheduling. Unlike uniform supervision [6] and task-based weighting [31], our hierarchical weighting $w_{t}=\lambda^{t-1}$ accounts for gradient magnitude decay specific to recursive architectures. CGAR achieves $1.71\times$ training speedup while maintaining competitive accuracy, with ablations revealing progressive depth curriculum alone provides $2.26\times$ speedup, demonstrating benefits beyond computational savings.
@@ III Method
Our goal is to accelerate the training speed of Tiny Recursive Model (TRM) [6], which suffers from training inefficiency due to fixed-depth training at all epochs (the original TRM paper reports $\sim\!36$ GPU-hours per dataset on their hardware). We propose CGAR (Curriculum-Guided Adaptive Recursion), a training methodology that achieves $1.71\times$ speedup through two complementary techniques: progressive depth curriculum (PDC) and hierarchical supervision weighting (HSW). Under controlled conditions on A100 GPU, CGAR reduces training time from $10.93$ hours (our replicated TRM baseline) to $6.38$ hours. We begin by reviewing TRM’s architecture and identifying its training inefficiencies, then present our solutions with mathematical formulations and theoretical analysis.
@@ III-A Background: TRM Architecture and Training
TRM trains through deep supervision [8], attaching losses to all $N_{\text{sup}}$ steps with uniform weighting $w_{t}=1/N_{\text{sup}}$. Let $h_{\text{out}}:\mathbb{R}^{L\times D}\rightarrow\mathbb{R}^{L\times V}$ denote the output projection head. The training objective is:
@@
Figure 2: CGAR Training with Progressive Curriculum and Hierarchical Weighting
@@ IV Experiments
Following TRM [6], we report exact match accuracy $\text{Acc}_{\text{exact}}:=N^{-1}\sum_{i=1}^{N}\mathds{1}[\hat{\bm{y}}_{i}=\bm{y}_{i}^{*}]$ measuring puzzles with all $81$ tokens correct (primary metric, since partial solutions violate global constraints), and token accuracy $\text{Acc}_{\text{token}}:=(NL)^{-1}\sum_{i=1}^{N}\sum_{j=1}^{L}\mathds{1}[\hat{y}_{ij}=y_{ij}^{*}]$ measuring per-position correctness (auxiliary metric). We evaluate CGAR on Sudoku-Extreme to answer four key research questions:
@@ IV-A Model Configuration
We implement CGAR using TinyRecursiveReasoningModel architecture with hidden dimension $D=512$, transformer blocks having $n_{L}=2$ layers with $h=8$ attention heads and feed-forward dimension $D_{\text{FFN}}=2048$, recursion depth $(n,T)$ following curriculum $\mathcal{C}_{\text{PDC}}(\rho)$ from Section III, deep supervision at $N_{\text{sup}}=16$ steps with hierarchical weighting $w_{t}=\lambda^{t-1}/Z_{\lambda}$ where $\lambda=0.7$ and $Z_{0.7}=3.283$, and total parameters $p\approx 5.0\text{M}$ identical to baseline TRM. The three-stage curriculum transitions at progress $\rho\in\{0.3,0.6,1.0\}$ with depths $(n,T)=(2,1)$ giving $\mathcal{D}_{\text{eff}}=6$ for $\rho<0.3$, then $(4,2)$ with $\mathcal{D}_{\text{eff}}=20$ for $0.3\leq\rho<0.6$, and $(6,3)$ with $\mathcal{D}_{\text{eff}}=42$ for $\rho\geq 0.6$. For controlled comparison, we train a baseline TRM on identical hardware (A100 GPU) using fixed $(n,T)=(6,3)$ with uniform weighting $w_{t}=1/16$ across all epochs following the TRM protocol [6].
@@ IV-B Training Configuration
We optimize with AdamW [39] using learning rate $\eta=5\times 10^{-4}$ with cosine annealing linear warmup [40] over 1,0001{,}000 steps ($\sim\!1.5\%$ of training), weight decay $\lambda_{\text{wd}}=0.01$, batch size $B=768$ on single GPU, gradient clipping at max norm $1.0$, and FP16 mixed precision with automatic loss scaling. Training runs for E=50, [^0] E=50{,}000 epochs following TRM protocol, where each epoch processes one minibatch of $B$ samples yielding $\sim\!65{,}000$ total optimization steps. We save checkpoints every 5,0005{,}000 epochs for progressive evaluation. All experiments use single NVIDIA A100 GPU (80GB VRAM) with CUDA 11.8 and PyTorch 2.0. Under these controlled conditions, CGAR training completes in $6.38$ hours wall-clock time versus $10.93$ hours for our replicated baseline TRM, achieving $1.71\times$ speedup.
@@ IV-C Implementation Details
Following TRM [6], we detach gradients for first T1T-1 H-cycles (Algorithm 2 lines 11-15), backpropagating only through final cycle $j=T$, preventing gradient explosion while maintaining supervision at all $N_{\text{sup}}$ steps. We implement ACT [9] with learned halting head $h_{\text{halt}}:\mathbb{R}^{L\times D}\rightarrow[0,1]$ predicting probabilities $q_{t}\in[0,1]$ supervised via binary cross-entropy $\ell_{\text{BCE}}(q_{t},\mathds{1}[\hat{\bm{y}}_{t}=\bm{y}^{*}])$ during training, halting at inference when $q_{t}>0.5$ for all positions. Curriculum progress $\rho_{e}=e/E$ computes at epoch start, determining depth $(n_{e},T_{e})=\mathcal{C}_{\text{PDC}}(\rho_{e})$ applied uniformly across all samples that epoch, ensuring stable gradient statistics within optimization steps. For reproducibility, we set random seed $42$ for PyTorch and NumPy.
@@ IV-D Evaluation Protocol
During training, we log primary loss $\mathcal{L}_{\text{CGAR}}$ from Eq. (6), training exact and token accuracies on current minibatch, curriculum progress $\rho$ with current depth $(n,T)$, ACT halting accuracy and loss, and gradient norms $\|\nabla_{\theta}\mathcal{L}\|_{2}$ every $100$ optimization steps. After training completion, we evaluate all $5$ saved checkpoints (epochs 30,00030{,}000 to 50,00050{,}000 at 5,0005{,}000 intervals) on full held-out test set ($N_{\text{test}}=200{,}000$ puzzles), computing exact and token accuracies, measuring inference speed (puzzles/second), analyzing halting step distributions $\{\delta(\bm{x}):\bm{x}\in\mathcal{D}_{\text{test}}\}$, and calculating generalization gap $\mathcal{R}(\theta):=\text{Acc}_{\text{test}}-\text{Acc}_{\text{train}}$. We compare CGAR against our replicated TRM baseline trained on identical hardware (A100 GPU) under the same experimental conditions. Note that the original TRM paper [6] reports $87.4\%$ exact accuracy with $\sim\!36$ hours training on different hardware, but we conduct our primary comparison using our controlled baseline for fairness.